OpenStudy (df001):

How do you know if a sequence is a geometric sequence? Please explain. I am trying to find r for this sequence and whether it is or not a geometric sequence. 1/2, 2, 8, 32, ...

OpenStudy (df001):

help

Nnesha (nnesha):

\[\huge\rm r= \frac{ a_2 }{ a_1 } , \frac{ a_4}{ a_3 }\] to find ratio divide next term by previous one you a_4/ a_3 = ? a_2 /a_1 = ?

OpenStudy (df001):

wait how do i know if the given sequence is geometric?

OpenStudy (df001):

my teacher said times by common factor i think

Nnesha (nnesha):

yea first divide terms if you get same number then that's mean sequence is increasing by same value so it's geometric if u get different number then it's not a geometric sequence

OpenStudy (df001):

what or how do I know what to divide the terms by?

OpenStudy (df001):

Am I suppose to divide by the previous terms

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm r= \frac{ a_2 }{ a_1 } , \frac{ a_4}{ a_3 }\] to find ratio divide next term by previous one you a_4/ a_3 = ? a_2 /a_1 = ? \(\color{blue}{\text{End of Quote}}\) here a represent terms a_2 = 2nd term a_3 = 3rd term so on

Nnesha (nnesha):

yes right

OpenStudy (df001):

ohh, common ratio! I get it now

OpenStudy (df001):

thanks

Nnesha (nnesha):

yep right so it's geometric or nope ?

OpenStudy (df001):

yes

OpenStudy (df001):

common ratio is 4, in other words, ratio is 4

Nnesha (nnesha):

yes right so if you divide a_4 by 3rd term you would get same answer which is 4 so it's a geometric sequence